OpenStudy (anonymous):

Rewrite the expression log5 (2√(x +1)3 / 3√3x + 7) in a form with no logarithms of products, quotients, or powers.

OpenStudy (anonymous):

\[\log base 5(2\sqrt{(x+1)_{}^{3}} / \sqrt[3]{3x+7})\]

OpenStudy (anonymous):

\[\log_5({2\sqrt{(x+1)^3} \over 3\sqrt{3x+7}})\] Is this the expression?

OpenStudy (anonymous):

Yes, that's it :)

OpenStudy (anonymous):

The second one is \[^{\sqrt[3]{3x +7}}\]

OpenStudy (anonymous):

OMG I didn't say that! Just a minute, I have to do it again.

OpenStudy (anonymous):

It's okay, i will wait a while :)

OpenStudy (anonymous):

So is the numerator \[\sqrt[2]{(x+1)^3}?\]

OpenStudy (anonymous):

\[ 2\sqrt{(x+1)^{3}} \over \sqrt[3]{3x+7}\]

OpenStudy (anonymous):

better?

OpenStudy (anonymous):

with the \[\log _{5}\] infront

OpenStudy (anonymous):

Good, that's much better :D \[\log_5{2\sqrt{(x+1)^3} \over \sqrt[3]{3x+7}}=\log_5(2)+\log_5(x+1)^{3 \over 2}-\log_5(3x+7)^{1 \over 3}\] \[=\log_5(2)+{3 \over 2}\log_5(x+1)-{1 \over 3}\log_5(3x+7)\]

OpenStudy (anonymous):

Ah..thank you so much, I really appreciate it!!!I get more help online than in my own country...You are a life-saver! :)